OpenStudy (anonymous):

Simplify the expression –3(x + 3)2 – 3 + 3x. What is the simplified expression in standard form? –3x2 – 18x – 27 –3 – 15x – 30 –3 + 3x + 6 –3 + 3x – 30

OpenStudy (anonymous):

Chu guava!!!

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[-3(x+2)^2-3+3x\] right?

OpenStudy (anonymous):

I keep getting 3x + 3 + 2

OpenStudy (welshfella):

-3 ( x + 3)^2 - 3 + 3x = -3(x^2 + 6x + 9) - 3 + 3x = -3x^2 - 18x - 27 - 3 + 3x

OpenStudy (welshfella):

can you continue - just add like terms

OpenStudy (welshfella):

I'm assuming that the 2 after the parentheses means 'squared'